### v1

# TDRI: Two-Phase Dialogue Refinement and Co-Adaptation for Interactive Image Generation

Abstract
Although text-to-image generation technologies have made significant advancements, they still face challenges when dealing with ambiguous prompts and aligning outputs with user intent.Our proposed framework, TDRI (Two-Phase Dialogue Refinement and Co-Adaptation), addresses these issues by enhancing image generation through iterative user interaction. It consists of two phases: the Initial Generation Phase, which creates base images based on user prompts, and the Interactive Refinement Phase, which integrates user feedback through three key modules. The Dialogue-to-Prompt (D2P) module ensures that user feedback is effectively transformed into actionable prompts, which improves the alignment between user intent and model input. By evaluating generated outputs against user expectations, the Feedback-Reflection (FR) module identifies discrepancies and facilitates improvements. In an effort to ensure consistently high-quality results, the Adaptive Optimization (AO) module fine-tunes the generation process by balancing user preferences and maintaining prompt fidelity. Experimental results show that TDRI outperforms existing methods by achieving 33.6% human preference, compared to 6.2% for GPT-4 augmentation, and the highest CLIP and BLIP alignment scores (0.338 and 0.336, respectively). In iterative feedback tasks, user satisfaction increased to 88% after 8 rounds, with diminishing returns beyond 6 rounds. Furthermore, TDRI has been found to reduce the number of iterations and improve personalization in the creation of fashion products. TDRI exhibits a strong potential for a wide range of applications in the creative and industrial domains, as it streamlines the creative process and improves alignment with user preferences.
keywords:
Controllable Low-light Enhance , Large Language Model , Diffusion Model , Prompt-Driven segmentation
journal: Neurocomputing11footnotetext: These authors contributed equally to this work.
[…]
[XDU] organization=Xidian University, addressline=No. 2 Taibai South Road, city=Xi’an, postcode=710071, state=Shaanxi, country=China
[…]
[UESTC] organization=University of Electronic Science and Technology of China, addressline=Qingshuihe Campus, 2006 Xiyuan Ave, West Hi-Tech Zone, city=Chengdu, postcode=611731, state=Sichuan, country=China
[…]
[XMU] organization=Xiamen University Malaysia, addressline=Jalan Sunsuria, Bandar Sunsuria, city=Sepang, postcode=43900, state=Selangor, country=Malaysia
[…]
[PKU] organization=Peking University, addressline=5 Yiheyuan Road, Haidian District, city=Beijing, postcode=100871, state=Beijing, country=China
[…]
[UofT] organization=Faculty of Applied Science & Engineering, University of Toronto, addressline=27 King’s College Cir, city=Toronto, postcode=M5S 1A1, state=Ontario, country=Canada
[…]
[SIGS] organization=Shenzhen International Graduate School, Tsinghua University, addressline=University Town of Shenzhen, Nanshan District, city=Shenzhen, postcode=518055, state=Guangdong, country=China
Figure 1: A multi-round dialogue interaction where the user refines the parrot’s appearance using the Dialogue-to-Prompt (D2P) module. The system updates the image based on user feedback and pose constraints.
1 Introduction
Generative artificial intelligence (AI) has made major strides in transforming industries through the automation of creative and non-creative tasks, which has contributed substantially to economic growth. Recent advancements in models like DALL·E 3 [1], Imagen [2], Stable Diffusion [3], and Cogview 3 [4] have revolutionized the field of image generation, enabling the creation of unique, high-quality images from textual descriptions [5]. Using these models, users are able to transform textual concepts into visually appealing and realistic images, sparking innovation across a variety of industries. Nevertheless, there is still considerable room for improvement, especially in terms of improving image resolution, precision, and capturing the subtleties of the input text. Furthermore, many existing systems still lack sufficiently intuitive user interfaces, which hampers their accessibility and interaction with users  [6].
Today, one of the greatest challenges in generative AI is the inability of these models to fully comprehend and accurately capture the subtle intentions behind user inputs. Although these models are capable of generating visually appealing images, they often fail to align with users’ exact mental models, particularly when textual prompts are complex or require fine-grained details  [7]. A major contributor to this issue is the intricate nature of human intent, where even small changes in words can have a significant impact on the generated results [8, 9]. Consequently, there is a persistent gap between user expectations and the results generated by AI system, often resulting in multiple iterations before the desired outcome is obtained  [10]. A frequent problem encountered by users is the need to repeat adjustments to inputs, resulting in a time consuming and inefficien process [11, 12].
Additionally, non-expert users may have difficulty manipulating input variables such as style, composition, and perspective, since they lack the necessary knowledge of prompt engineering to control these elements effectively [4]. Due to this complexity, as well as the inherent variability of the generated images, users have often been forced to engage in a trial-and-error process in order to achieve satisfactory results [13]. For example, identical prompts may result in significantly different outputs in terms of layout, color scheme, and perspective, requiring users to modify their inputs numerous times until the final result is consistent with their vision [8, 9, 14].
To address these challenges, we introduce the TDRI (Text-driven Iterative Refinement Interaction) framework, which is designed to enhance the user experience by refining image outputs through iterative feedback. By contrast to traditional methods that rely heavily on prompt engineering, TDRI allows users to interact with the model through multiple rounds of feedback, gradually refining the generated image. Through this continuous feedback loop, the model is able to more accurately capture the user’s intent, thereby reducing the number of trial-and-error attempts and improving the efficiency of the process [4, 13]. By maintaining this iterative process, TDRI simplifies the image generation workflow, making it more user-friendly and accessible to both novice and expert users alike.
The key innovation of the TDRI framework is its ability to combine external user inputs with internal optimization processes, offering a seamless, intuitive human-machine interaction. The dialogue-based iterative approach reduces the complexity of input operations and streamlines the image generation process, which results in a more accurate and efficient image generation process. With this approach, TDRI facilitates better alignment between user expectations and the generated outputs, thereby significantly increasing the quality and relevance of the resulting images.
Our main contributions are as follows:
  • 1.
    Our study examines specialized human-machine interaction techniques for guiding users through the refined image generation process in order to ensure their intentions are accurately captured and reflected in the final images.
  • 2.
    The TDRI framework, a two-phase dialogue-based methodology for interactive image generation, is introduced, combining external user interactions with internal optimization processes to ensure that textual concepts are accurately translated into visual outputs.
  • 3.
    We demonstrate the applicability of TDRI across various image generation tasks, showcasing its versatility in rapidly visualizing and iterating on a diverse range of creative concepts.
2 Related work
Artificial intelligence continues to demonstrate groundbreaking progress across interdisciplinary fields, spanning foundational vision technologies [15, 16, 17, 18, 19], cognitive visual systems [20, 21, 22], and intelligent engineering solutions [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. This review specifically examines transformative breakthroughs in generative AI, focusing on image synthesis innovations [36, 37, 38, 39] that redefine content creation paradigms.
Text-Driven Image Editing Framework
Recent advancements in text-to-image generation have focused on aligning models with human preferences, using feedback to refine image generation. Studies range from Hertz et al. [40]’s framework, which leverages diffusion models’ cross-attention layers for high-quality, prompt-driven image modifications, to innovative methods like ImageReward [41], which develops a reward model based on human preferences. These approaches collect rich human feedback [42, 43], from detailed actionable insights to preference-driven data, training models for better image-text alignment and adaptability [44] to diverse preferences, marking significant progress in personalized image creation.
Ambiguity Resolution in Text-to-Image Generation
From visual annotations [45] and model evaluation benchmarks [46] to auto-regressive models [47] for rich visuals, along with frameworks for abstract [48] and inclusive imagery [49], the text-to-image field is advancing through strategies like masked transformers [50], layout guidance [51] without human input, and feedback mechanisms [43] for quality. The TIED framework and TAB dataset [52] notably enhance prompt clarity through user interaction, improving image alignment with user intentions, thereby boosting precision and creativity.
Human Preference-Driven Optimization for Text-to-Image Generation Models
Zhong et al. [53] significantly advance the adaptability of LLMs to human preferences with their innovative contributions. Zhong et al.’s method stands out by leveraging advanced mathematical techniques for a nuanced, preference-sensitive model adjustment, eliminating the exhaustive need for model retraining. Xu et al. [41] takes a unique approach by harnessing vast amounts of expert insights to sculpt their ImageReward system, setting a new benchmark in the creation of images that resonate more deeply with human desires. Together, these advancements mark a pivotal shift towards more intuitive, user-centric LLMs technologies, heralding a future where AI seamlessly aligns with the complex mosaic of individual human expectations.
Figure 2: An overview of the two-phase framework TDRI. (a) In the Initial Generation Phase, the system processes user prompts via a U-Net-based diffusion model, generating base images with pose constraints. (b) In the Interactive Refinement Phase, user feedback is integrated to iteratively refine the image through dialogue-to-prompt generation, ambiguity scoring, and adaptive optimization.
3 Proposed method
We propose a two-phase framework for image generation in multi-turn dialogues: the Initial Generation Phase, where the system processes the user’s initial prompt ($w_{1}$) to generate an image ($I_{1}$) and extract pose ($pose_{1}$) as a constraint, and the Interactive Refinement Phase, where three modules—Dialogue-to-Prompt ($D2P$), Feedback-Reflection ($F_{R}$), and Adaptive Optimization ($A_{O}$)—iteratively refine the image based on user feedback to ensure comprehensive prompt representation.
3.1 Initial Generation Phase
The Initial Generation Phase initializes the image generation by processing the user input prompt $w_{1}$. The system generates a base image $I_{1}$ using a prompt-conditioned generative model $G(\cdot)$: $I_{1}=G(w_{1})$, where $I_{1}$ is the initial image generated based on prompt $w_{1}$. Subsequently, a pose estimator $\mathcal{P}(\cdot)$ extracts the pose $pose_{1}$ from $I_{1}$, represented by keypoint coordinates $\{(x_{i},y_{i})\}_{i=1}^{K}$ for $K$ keypoints: $pose_{1}=\mathcal{P}(I_{1})$. The extracted pose $pose_{1}$ acts as a structural constraint for subsequent iterations. A Gaussian smoothing function $\mathcal{S}(\cdot)$ is applied to refine $pose_{1}$, expanding its influence: $pose_{1}^{\prime}=\mathcal{S}(pose_{1})$. This refined pose $pose_{1}^{\prime}$ is used as a guiding feature in future image generation rounds, maintaining core structural integrity while allowing flexibility in user-directed updates.
3.2 Interactive Refinement Phase
3.2.1 Dialogue-to-Prompt Module ($D2P$)
The Dialogue-to-Prompt Module ($D2P$) formulates the prompt $P_{t}$ at each timestep $t$ by integrating the dialogue history $h_{t}$ and the latest user input $w_{t}$. The dialogue history is defined as:
$h_{t}=\{(w_{1},r_{1}),(w_{2},r_{2}),\dots,(w_{t-1},r_{t-1})\},$ (1)
where $w_{i}$ and $r_{i}$ represent the user input and system response at step $i$, respectively. The Summarizer $M_{S}$ synthesizes $h_{t}$ and $w_{t}$ to generate $P_{t}$:
$\begin{split}P_{t}&=M_{S}(h_{t},w_{t})\\ &=g_{\text{sum}}\left(\sum_{i=1}^{t-1}\lambda_{i}\phi(w_{i})+\mu_{i}\psi(r_{i}),\phi(w_{t})\right),\end{split}$ (2)
where $\lambda_{i}$, $\mu_{i}$ are weighting coefficients, $\phi(\cdot)$, $\psi(\cdot)$ are embedding functions mapping inputs to high-dimensional feature spaces, and $g_{\text{sum}}$ denotes the summarization operation. This aggregation ensures that $P_{t}$ encapsulates both historical context and current user intent, optimizing it for image generation. Subsequently, the Generation Model $M_{G}$ utilizes $P_{t}$ to produce the image $I_{t}$, conditioned on the initial pose $pose_{1}^{\prime}$ and accumulated context $\mathcal{C}_{t-1}$:
$I_{t}=M_{G}(P_{t}\mid pose_{1}^{\prime},\mathcal{C}_{t-1}),$ (3)
where $\mathcal{C}_{t-1}$ aggregates contextual information from prior iterations. This iterative update mechanism enables dynamic adaptation to user feedback, refining $I_{t}$ to align with evolving user preferences across multiple dialogue turns.
3.2.2 Feedback-Reflection Module ($F_{R}$)
The Feedback-Reflection Module ($F_{R}$) evaluates the generated image $I_{t}$ by extracting a set of descriptive features or captions, $C_{t}=\{C_{t}^{1},C_{t}^{2},\dots,C_{t}^{N}\}$, where each $C_{t}^{i}$ represents a distinct characteristic of the image. In our implementation, the extraction function $f_{E}$ is handled by a vision-language model (VLM), specifically Qwen-VL[54]. We incorporate specific prompt templates to guide the VLM in assessing the completeness of the generated image, prompting it to identify essential elements, such as objects, colors, and other critical features:
$C_{t}=f_{E}(I_{t})=\left\{C_{t}^{i}\mid i=1,2,\dots,N\right\},$ (4)
where $f_{E}$ maps the image $I_{t}$ to a structured description $C_{t}$. The extracted features $C_{t}$ provide a comprehensive evaluation of the image, which is then compared to the input prompt $P_{t}$ to assess how well the image aligns with user expectations and identify areas for further refinement.
To evaluate the consistency between $P_{t}$ and $C_{t}$, a similarity measure $\sigma(P_{t},C_{t})$ is used to compute the discrepancy between the prompt and generated image. This results in an ambiguity score $r_{t}$: $r_{t}=1-\sigma(P_{t},C_{t})$, where $r_{t}\in[0,1]$ indicates the level of mismatch. The function $\sigma(P_{t},C_{t})$ is defined as:
$\sigma(P_{t},C_{t})=\frac{\sum_{i=1}^{N}\nu_{i}\kappa(P_{t}^{i},C_{t}^{i})}{\sum_{i=1}^{N}\nu_{i}},$ (5)
where $\kappa(P_{t}^{i},C_{t}^{i})$ represents a similarity function between the $i$-th component of the prompt and the corresponding feature in the generated image, and $\nu_{i}$ denotes a weight assigned to each feature’s importance in the evaluation.
When the ambiguity score $r_{t}$ exceeds a threshold $\tau$, the system seeks further user input to refine the prompt. This process generates a clarification query $q_{t+1}$, which is formulated as:
$q_{t+1}=f_{\text{clarify}}(P_{t},C_{t},r_{t}),$ (6)
where $f_{\text{clarify}}$ is a function that analyzes the prompt $P_{t}$, image captions $C_{t}$, and the ambiguity score $r_{t}$ to determine the most relevant aspect of the ambiguity. It then constructs a clarification query accordingly, targeting the part of the input that requires further refinement. By iteratively calculating $r_{t}$ and generating $q_{t+1}$, the system continuously aligns its output with the user’s evolving intent, optimizing the prompt $P_{t}$ and the resulting image $I_{t}$ over multiple dialogue rounds.
3.2.3 Adaptive Optimization Module ($A_{O}$)
The Adaptive Optimization Module ($A_{O}$) integrates Direct Preference Optimization ($DPO$) and Attend-and-Excite ($A\&E$) to ensure alignment between generated images and user preferences while maintaining prompt fidelity.
Direct Preference Optimization ($DPO$) leverages user preference pairs $\mathcal{P}=\{(x_{w},x_{l})\}$, where $x_{w}$ is the preferred image and $x_{l}$ is the less preferred one. The goal is to maximize the likelihood of generating $x_{w}$ over $x_{l}$, which is formalized as:
$\mathcal{L}_{DPO}(\theta)=\mathbb{E}_{(x_{w},x_{l})\sim\mathcal{P}}\left[\log\frac{\pi_{\theta}(x_{w}\mid s)}{\pi_{\theta}(x_{l}\mid s)}\right].$ (7)
Attend-and-Excite ($A\&E$) ensures that all key elements from the input prompt $P_{t}$ are adequately represented in the image $I_{t}$. The misalignment loss is defined as:
$L=1-Sim(I_{t},P_{t}),$ (8)
where the similarity score $Sim(I_{t},P_{t})$ measures the alignment between the image and the prompt. The gradient $\Delta P_{t}=\nabla_{P_{t}}L$ is computed to identify under-represented elements, which are then used to refine the prompt and regenerate the image.
During training, ControlNet is tuned using the combined loss function:
$\mathcal{L}_{A_{O}}(\theta)=\mathcal{L}_{DPO}(\theta)+\lambda\mathcal{L}_{A\&E}(\theta),$ (9)
where $\lambda$ controls the balance between preference alignment and prompt fidelity.
4 Experiment
We evaluated the performance of the TDRI framework in two scenarios: fashion product creation and general image generation. Each scenario presents unique requirements. We first focused on fashion product creation due to the availability of a larger dataset, allowing us to capture fine-grained intent and user preferences. After demonstrating the model’s success in this domain, we extended the framework to the general image generation task, where the focus shifted towards satisfying broader user intent.
4.1 Q&A Software Annotation Interface
Image Panel: Two images are displayed side-by-side for comparison or annotation. These images seem to depict artistic or natural scenes, suggesting the software can handle complex visual content. HTML Code Snippet: Below the images, there’s an HTML code snippet visible. This could be used to embed or manage the images within web pages or for similar digital contexts. Interactive Command Area: On the right, there is an area with various controls and settings: Current task and image details: Displayed at the top, indicating the task at hand might be related to outdoor scenes. Navigation buttons: For loading new images and navigating through tasks. Annotation tools: Options to add text, tags, or other markers to the images. Save and manage changes: Buttons to save the current work and manage the task details.
Figure 3: Screenshot of the Q&A software annotation interface.
Objective Accurately describe and tag visual content in images to train our machine learning models.
Steps
  1. 1.
    Load Image: Use the ’Load Image’ button to begin your task.
  2. 2.
    Analyze and Describe:
    • (a)
      Examine each image for key features.
    • (b)
      Enter descriptions in the text box below each image.
  3. 3.
    Tagging:
    • (a)
      Apply relevant tags from the provided list to specific elements within the image.
  4. 4.
    Save Work: Click ’Save Task’ to submit your annotations. Use ’Load Last’ to review past work.
Guidelines
  • 1.
    Accuracy: Only describe visible elements.
  • 2.
    Consistency: Use the same terms consistently for the same objects or features.
  • 3.
    Clarity: Keep descriptions clear and to the point.
Support For help, access the ’Help’ section or contact the project manager at [contact information].
Note: Submissions will be checked for quality; maintain high standards to ensure data integrity.
4.2 Task 1 : Fashion Product Creation
Figure 4: This image presents a variety of fashion models and outfits, segmented by user preferences, showcasing styles from elegant dresses to casual and professional jackets, modeled by individuals of diverse ethnicities.
Figure 5: Human Voting for Statement: Direct Preference Optimization can improve generation results.
4.2.1 Setting
Fashion product creation poses greater challenges than general image generation due to higher demands for quality and diversity. Our Agent system requires advanced reasoning and multimodal understanding, supported by ChatGPT-4 for reasoning tasks. For image generation, we used the SD-XL 1.0 model, fine-tuned with the DeepFashion dataset [55] for clothing types and attributes. The LoRA [56] method was applied for fine-tuning on four Nvidia A6000 GPUs, resulting in more consistent outputs. To provide a personalized experience, we trained multiple models with different ethnic data, allowing users to choose according to preferences. Using Direct Preference Optimization (DPO), model parameters were updated after every 40 user feedback instances, repeated three times, with the DDIM sampler for image generation.
4.2.2 Result Analysis
Figure 4 showcases the outputs of six models optimized through Direct Preference Optimization (DPO) based on feedback from six users. Each model generated fashion products using the same prompt and random seed, with variations reflecting individual user preferences. The rows represent the six users, while the columns display different outfit types, including sheer sleeve dresses, floral casual dresses, sweater skirts, winter jackets, and formal suits. Each row is divided into base model outputs, trained on general user group characteristics (e.g., "Asian people Model" or "Black people Model"), and DPO-tuned outputs, personalized using user-specific interaction data. The results highlight how DPO influences the latent space to produce tailored outputs, even with identical prompts and random seeds, effectively aligning with diverse user preferences.
Figure 5 illustrates the results of human evaluations on the effectiveness of Direct Preference Optimization (DPO) in improving generation results. Each chart corresponds to feedback from a specific user (User1 to User6) and represents their voting distribution across five levels: "Strongly Disagree," "Neutral," and "Strongly Agree," with a purple arrow indicating the median response. The bar heights reflect the number of topics rated at each level. Most users (Users 1 through 6) showed a strong preference for DPO-optimized outputs, as indicated by the majority of votes falling into the "Agree" or "Strongly Agree" categories. The median responses consistently lean toward positive agreement, highlighting significant performance improvements achieved through DPO fine-tuning.
4.3 Task 2: General Image Generation
Figure 6: Comparison of cherry blossom tea images generated across four rounds by various models.
Figure 7: Iterative refinement process of image generation across four rounds for various topics using TDRI. Each row represents a specific topic, showing progressive improvements in alignment with user intent.
4.3.1 Setting
In this task, the Summarizer generates prompts by aggregating the user’s input, which are then used to create images. These images are captioned by Qwen-VL [54], a Vision-Language Model, across seven aspects: ’Content’, ’Style’, ’Background’, ’Size’, ’Color’, ’Perspective’, and ’Others’. We compare the CLIP similarity scores between the current generated image and each caption to identify ambiguous aspects. One of the three lowest-scoring aspects is randomly selected for questioning, and the user can choose to respond. In human-in-the-loop image generation, a target reference image is set, and user feedback is provided after each generation, with similarity to the target image used to assess effectiveness.
4.3.2 Data Collection
We curated 496 high-quality image-text pairs from the ImageReward dataset [41], focusing on samples with strong alignment to prompts. By removing abstract or overly complex prompts, as very long prompts tend to reduce accuracy and fail to clearly reflect the user’s intent, we included people, animals, scenes, and artworks. Over 2000 user-generated prompts were used, with some images containing content not explicitly mentioned in the prompts. Each sample underwent at least four dialogue rounds for generation.
Figure 8: Human Voting for Statement: Multi-turn dialogues can approximate the user’s potential intents.
Table 1: Evaluations of prompt-intent alignment, image-intent alignment, and human voting across various methodologies and integrations. Augmentation refers to using LLMs to infer ambiguity and enhance the initial prompt. TDRI-Reflection is the interaction refinement phase of our TDRI.
Methods Prompt-Intent Alignment Image-Intent Alignment Human Voting
T2I CLIPscore T2I BLIPscore I2I CLIPscore I2I BLIPscore
GPT-3.5 augmentation 0.154 0.146 0.623 0.634 5%
GPT-4 augmentation 0.162 0.151 0.647 0.638 6.2%
LLaMA-2 augmentation 0.116 0.133 0.591 0.570 6.1%
Yi-34B augmentation 0.103 0.124 0.586 0.562 4.3%
TDRI-Reflection 0.281 0.285 0.753 0.767 25.8%
TDRI-Reflection + ImageReward RL [41] 0.297 0.284 0.786 0.776 26.5%
[…]
4.3.3 Baseline setup
To demonstrate the effectiveness of our Reflective Human-Machine Co-adaptation Strategy in uncovering users’ intentions, we established several baselines. One method to resolve ambiguity in prompts is using Large Language Models (LLMs) to rewrite them. We employed various LLMs, including ChatGPT-3.5, ChatGPT-4 [57], LLaMA-2 [58], and Yi-34B [59].
The table 1 evaluates methods for aligning generated prompts with target intents and images, using metrics like T21 CLIPscore, T21 BLIPscore, and Human Voting. Compared methods include augmentation techniques (e.g., GPT-3.5, GPT-4, LLaMA-2) and TDRI (ours) with iterative refinement. TDRI (ours) outperforms all other methods, achieving the highest scores: 0.338 in Prompt-Intent Alignment, 0.812 in Image-Intent Alignment, and 33.6% in Human Voting. Augmentation methods performed poorly, with human voting results between 4.3% and 6.2%, while TDRI-Reflection variants improved results to 25.8% and 26.5%. In conclusion, TDRI demonstrates clear superiority in generating outputs aligned with target intents, highlighting the effectiveness of its optimization approach. Experiments using SD-1.4 with the DDIM sampler on Nvidia A6000 GPUs confirm its high performance in generative tasks.
Figure 6 shows a comparison of cherry blossom tea images generated across four iterative rounds by different models, including TDRI (Ours), CogView 3, Imagen 3, PTP, and DALL·E 3. Each round introduces additional refinements to the prompt, showcasing the models’ abilities to adapt to specific details such as ’wooden table,’ ’top view,’ and ’flowers float on tea. Figure 7 illustrates iterative the refinement process of image generation across four rounds using TDRI, showcasing its ability to progressively align outputs with user intent. Each row represents a distinct topic—cherry blossom tea, parrot, teenage girl, and Asian temple—demonstrating enhanced detail and accuracy through incremental prompt refinements. Figure 8 collects the approval ratings from five testers. In these dialogues, we explore whether the users agree that the multi-round dialogue format can approximate the underlying generative target. In most cases, HM-Reflection produces results closely aligned with user intent.
4.3.4 Qualitative Results
Embedding Refinement by Round: The t-SNE visualization in Figure 9 highlights how embeddings evolve across three interaction rounds. With each round of feedback, the embedding distribution becomes increasingly compact. It indicates that the model progressively refines its understanding of user intent, as seen by the tighter clustering of similar samples and reduced overlap between rounds. These improvements demonstrate the model’s ability to capture user preferences more effectively through iterative optimization (refer to Tables 1 and 2).
Figure 9: t-SNE visualization of embeddings across three interaction rounds.
Figure 10: Heatmap showing user perception of the model’s ability to capture intent across different dialogue rounds. The intensity peaks around 3 rounds.
User Perception of Intent Capture: Figure 10 presents a heatmap illustrating user perception of the model’s ability to capture intent across different dialogue rounds. The intensity peaks around the third round, indicating that users felt the model most accurately understood their intent at this stage. This suggests that by the third interaction, the model has significantly improved its comprehension of user preferences, and subsequent rounds provide only marginal gains in refining user intent.
User Interaction Distribution by Round: The distribution of user interactions across dialogue rounds is shown in Figure 11. The majority of users required around five rounds to refine their image generation, with the highest proportion (21.1%) achieving their desired results by the fifth round. This suggests that the TDRI framework effectively captures user preferences within a relatively small number of interactions, with diminishing returns in later rounds as fewer users required additional feedback beyond round five.
Figure 11: Proportion of users across dialogue rounds in the TDRI framework peaks at 5 rounds (21.1%), indicating most users refined their image generation within 5 interactions.
4.3.5 Quantitative Results
Table 2: Ablation study of multi-dialog models across different rounds and metrics (CLIP and BLIP scores).
Multi-dialog SD-1.4 SD-1.5 DALL-E 3 MetaGPT PTP CogView 3 Imagen 3
CLIP BLIP CLIP BLIP CLIP BLIP CLIP BLIP CLIP BLIP CLIP BLIP CLIP BLIP
Round 1 $0.728$ $0.703$ $0.723$ $0.699$ $0.651$ $0.674$ $0.646$ $0.672$ $0.661$ $0.681$ $0.643$ $0.664$ $0.671$ $0.691$
Round 2 $0.759$ $0.738$ $0.746$ $0.725$ $0.675$ $0.690$ $0.671$ $0.691$ $0.682$ $0.700$ $0.667$ $0.679$ $0.696$ $0.712$
Round 3 $0.776$ $0.764$ $0.773$ $0.784$ $0.691$ $0.718$ $0.689$ $0.711$ $0.701$ $0.716$ $0.684$ $0.696$ $0.727$ $0.732$
Round 4 $0.804$ $0.824$ $0.790$ $0.811$ $0.743$ $0.736$ $0.726$ $0.742$ $0.712$ $0.726$ $0.705$ $0.717$ $0.751$ $0.742$
Image Editing vs. From Scratch Generation As shown in Table 3, Image Editing significantly outperforms the From Scratch method in terms of consistency (0.88 vs. 0.75) and user satisfaction (90% vs. 78%). Additionally, Image Editing requires less time (9 minutes vs. 12 minutes). This indicates that editing an existing image rather than generating from scratch leads to a more refined and efficient process, aligning closely with user expectations.
Table 3: Effect of Interaction Turns on Image Quality, Satisfaction, and Time
Turns Satisfaction (%) CLIP Score Time (min)
2 70% 0.72 6
4 85% 0.78 9
6 87% 0.80 11
8 88% 0.81 12
Table 4: Comparison of Simple vs. Complex Prompts
Prompt Type Generation Success Rate (%) Average CLIP Score Human Voting (%)
Simple Prompts 92% 0.85 87%
Complex Prompts 65% 0.60 62%
Table 5: Generalized Model vs. Sample-Specific D3PO
D3PO Training Method User Satisfaction (%) Time to Convergence (iterations) CLIP Score
Generalized Model 83% 5 0.77
Sample-Specific Model 90% 8 0.85
Table 6: Attend-and-Excite Usage Frequency and T2I Similarity at Different Thresholds
Attend-and-Excite Threshold 0.80 0.75 0.73 0.70 0.68 0.66
Frequency of Usage 0 8.7 % 31.3 % 51.6 % 72.5 % 95.8 %
T2I Similarity Improvement 0 0.23 % 1.87 % 2.36 % 2.67 % 1.3 %
Complex Prompt Exclusion Justification Table 4 compares the performance of simple and complex prompts in generative tasks, highlighting significant differences in success rates and alignment with user intent. Simple prompts achieve a much higher Generation Success Rate (92%) compared to complex prompts (65%). Similarly, the Average CLIP Score is considerably better for simple prompts (0.85) than for complex prompts (0.60), indicating that simple prompts generate outputs more aligned with the intended target. In terms of Human Voting, simple prompts received a higher preference rate (87%) compared to complex prompts (62%). This decline in performance with complex prompts suggests that simplicity in prompts leads to more consistent and effective results. These findings support the focus on avoiding overly complex prompts to achieve better alignment and user satisfaction in generative models.
Generalized vs. Sample-Specific D3PO Table 5 compares the performance of the Generalized Model and the Sample-Specific Model in D3PO training. The Sample-Specific Model achieves higher User Satisfaction (90%) and a better CLIP Score (0.85) compared to the Generalized Model, which has a User Satisfaction rate of 83% and a CLIP Score of 0.77. However, the Sample-Specific Model requires more Iterations to Converge (8) than the Generalized Model (5), indicating higher computational costs. These results suggest that while sample-specific tuning produces higher-quality outputs and better aligns with user preferences, it comes at the expense of increased computation time. This trade-off highlights the need to balance performance and efficiency based on the specific requirements of a task.
Attend-and-Excite Performance: We also conducted independent experiments on Algorithm (Attend-and-Excite) using the dataset from Task 2. As shown in Table 6, the usage frequency of Attend-and-Excite varies with different thresholds $k$. At $k=0.72$ and $k=0.7$, the usage frequencies were 31.1% and 51.1%, respectively, with CLIP score increases of 1.8% and 2.3%, demonstrating that these settings improve image-text alignment.
Lightweight Models Comparison Table 7 compares the performance of lightweight models of different sizes (7B, 5B, and 3B) based on user satisfaction, CLIP score, and computation time. The 7B model achieves the highest User Satisfaction (90%) and CLIP Score (0.85), indicating superior image quality and alignment with user intent, but it requires the longest computation time (15 minutes). On the other hand, the 3B model is the fastest, with a computation time of just 6 minutes, but it compromises on performance, with a lower User Satisfaction (78%) and CLIP Score (0.77). The 5B model strikes a middle ground, offering improved performance over the 3B model with a User Satisfaction of 85% and a CLIP Score of 0.82 while reducing computation time to 10 minutes compared to the 7B model. These results highlight a trade-off between speed and performance, suggesting that the choice of model size should depend on the specific requirements of the task, balancing efficiency and quality.
Table 7: Performance Comparison of Lightweight Models
Model Size User Satisfaction (%) CLIP Score Computation Time (minutes)
7B 90% 0.85 15
5B 85% 0.82 10
3B 78% 0.77 6
5 Conclusion
This study introduced TDRI (Text-driven Iterative Refinement Interaction), a framework for interactive image generation that combines dialogue-driven interactions and optimization techniques to enhance personalization and alignment with user intent. Through its two-phase process—Initial Generation and Interactive Refinement—TDRI progressively improves outputs with user feedback, reducing trial-and-error and enhancing efficiency. Experiments demonstrated TDRI’s ability to deliver high-quality, personalized results across diverse tasks, outperforming existing methods in user satisfaction and alignment metrics. Its adaptability shows promise for applications in both creative and industrial domains. Future work will focus on addressing limitations, such as handling complex prompts, reducing computational costs, and integrating finer feedback mechanisms, to further optimize performance and broaden its applicability.
6 Limitations
While TDRI offers significant improvements, it has certain limitations. The model may struggle to accurately translate complex, multi-level prompts into images due to the VL model’s difficulty in capturing fine-grained details, leading to inaccurate captions. Additionally, cross-modal transfer errors can obscure user intent, reducing communication efficiency. The method is also computationally intensive and time-consuming, posing challenges for users with less powerful hardware. Future work should focus on enhancing efficiency and expanding the system’s ability to generalize across diverse inputs to improve real-world usability.